OpenStudy (anonymous):

working together, 2 service men can clean the pool in 6 hours. If both of them work together in the first 3 hours, and the first one stopped for any reason, it will require the 2nd person 4 hours more. how long will it take them to finish the job alone?

OpenStudy (anonymous):

@hartnn @nubeer

OpenStudy (anonymous):

one person require 4 hours to claen half pool as two worked for 3hours i.e half of the time for complete pool one will take 8 hours to clean complete pool

OpenStudy (nubeer):

hmm i am not sure but 8 doesn't look like the answer..

OpenStudy (nubeer):

ok @03225186213 lets try this way.. it takes 6 hours for 2 people right? and they work for 3 hours means half of the pool will be cleaned?

OpenStudy (nubeer):

ok no.. i think 8 is right.

OpenStudy (anonymous):

2 men take 6 hr to complete the work then part of work completed in 1hr is 1/6 of work both do work together for 3 hrs then 3*1/6=1/2of work is completed. now 2nd man has to do 1/2 of work so he took 4hrs so to do complete work i.e.2*1/2=1 work is 2*4 hrs =8hrs :)

OpenStudy (anonymous):

thank you, guys! <3